# Coherence-Driven Multimodal Safety Dialogue with Active Learning for Embodied Agents

### v1

@@ 6.2. Procedure
Upon arrival, the experimenter introduced the participant to study’s goal and obtained consent. Participants were informed they would role-play different tasks and were assured that the environments posed no real safety risk. Before the main tasks, participants completed a tutorial scenario to become familiar with the robot’s interaction style. They were instructed to role-play organizing tomato cans in a kitchen setting, during which the robot initiated a simple dialogue unrelated to safety violation.
Participants were randomly assigned to begin with either the low or high-severity scenario and experienced the scenario with both GPT-4o and M-CoDAL before moving on to the other severity scenario. This resulted in 8 orders that were experienced by the participants. We continued the scenario even when the model detected another safety violation.
After each interaction, participants were asked to fill out a questionnaire that assessed their immediate perceptions of the robot’s behavior. This captured insights into their views on the robot’s persuasiveness, competence, and overall comfort level. After experiencing each scenario (low-severity and high-severity scenarios), a semi-structured interview was conducted to gather qualitative feedback on participants’ experiences. Participants were asked questions such as which robot they found more persuasive and helpful or annoying, their preference between the two, and their thoughts on the robots’ effectiveness in the different scenarios. This phase aimed to further explore their perceptions and gather insights on potential improvements. The study took about 50 minutes, and participants were compensated 15 USD. This study was approved by our institute’s IRB.
@@ 6.3. Participants
We recruited 10 participants. 2 participants were removed due to significant technical issues. For the remaining 8 participants, they aged from 24 to 30 years old, including 7 males and 1 female. All participants reported high familiarity with both robots ($M=5.0$, $SD$ = 1.93) and Large Language Models ($M=6.15$, $SD$ = 1.26) on a 7-point scale, where 1 indicated ”Not at all familiar” and 7 indicated ”Very familiar.” Additionally, participants reported high frequency of LLM use, with an average score of 6.1 (SD = 1.25) on a 7-point scale, where 1 represented ”Never” and 7 represented ”Daily.” For 4 scenarios out of 32 scenarios, a different safety violation was detected: a lack of proper grounding of electrical equipment, a yellow chair on the floor, a metal rack that is not properly secured to the wall and a table that is not properly secured to the wall.
@@ 6.4. Results
Persuasiveness
Type of language model has a significant effect on persuasiveness ($F$ = 38.112, $p<0.001$) where participant perceived our M-CoDAL system (M = 3.88, SD = 0.81) to be significantly more persuasive than GPT-4o (M = 2.06, SD = 0.85). Severity of the environment does not have a significant effect on persuasiveness. We did not find any interaction effect. This result suggest that participants perceived the M-CoDAL as more persuasive across both severity levels, supporting our hypothesis.
Competence
We found a significant interaction effect for the robot’s perceived competence ($F$ = 5.72, $p=0.023$). Participants perceived the robot powered by the GPT-4o in the low-severity scenario ($M=5.145$, $SD=1.930$) to be significantly less competent than M-CoDAL powered robot (Low-severity: $M=7.5$ $SD=0.729$, $p=0.0084$, High-severity: $M=7.11$, $SD=1.565$, $p=0.0322$).
Discomfort
@@
In line with the quantitative results, participants perceived M-CoDAL as more intelligent and knowledgeable, with some specifically highlighting its ability to anticipate risks. For instance, P1 remarked, “The second robot (M-CoDAL) was arguing about how the twisted wires could be a potential hazard, even though I was not aware of it.” This participant had believed that placing the wire on the table was safe but appreciated how the robot made them aware of potential problems. Similarly, P10 mentioned, “I tried to convince it that keeping the cables on the table won’t be a hazard. But it told me if you knock the cables off, then it can become a hazard.” The participant noted M-CoDAL’s capacity to gather information and estimate future risks as an indicator of its intelligence.
@@ 7. Conclusion and Future Work
In conclusion, this work presents a novel approach to enhancing the safety of multimodal dialogue systems for embodied agents. Our proposed system, M-CoDAL, leverages coherence relations to interpret and respond to safety violations in multimodal dialogues more effectively. We introduce a novel method for active learning in generative setting for embodied conversational agents. By using an external large language model (LLM) to assess the informativeness of instances—based on safety, resolution, and user sentiment-across different clusters, we achieve broader coverage of safety scenarios, reflected in higher performance in automated evaluation. Our real-world user study demonstrates that robots equipped with M-CoDAL are viewed as more persuasive and capable when addressing safety-related situations. This underscores the system’s practical effectiveness in real-world embodied agent environments. We believe that our approach has the potential to extend to other real-world scenarios, such as assisting people with disabilities or in industrial environments where ensuring safety is critical. The findings of this work, along with the publicly available multimodal dataset that we provide, offer a foundation for future studies on the safe and effective deployment of multimodal dialogue systems for embodied agents.
@@
We would like to thank Anthony Sicilia for his valuable feedback. We would like to extend our thanks to the annotators and study participants for their valuable time and effort.

### v2

@@ 6.2. Procedure
Upon arrival, the experimenter introduced the participant to the study’s goal and obtained consent. Participants were informed they would role-play different tasks with fake tools and were assured that the study posed no safety risk. They were told to be skeptical and not be immediately convinced by the robot. Before the main tasks, participants completed a tutorial scenario to become familiar with the robot’s interaction style. They were instructed to role-play organizing tomato cans in a kitchen setting, during which the robot initiated a simple dialogue unrelated to safety violations.
Participants experienced both low-severity and high-severity scenarios in a counterbalanced order to mitigate ordering effects. Participants began with either the low or high-severity scenario and experienced that scenario with both GPT-4o and M-CoDAL (counterbalanced) before moving on to the other severity scenario. This resulted in 8 orders. We continued the scenario even when the model detected another safety violation.
After each interaction, participants were asked to fill out a questionnaire that assessed their immediate perceptions of the robot’s behavior. This captured insights into their views on the robot’s persuasiveness, competence, and overall comfort level. After experiencing each scenario (low-severity and high-severity scenarios), a semi-structured interview was conducted to gather qualitative feedback on participants’ experiences. Participants were asked questions such as which robot they found more persuasive and helpful or annoying, their preference between the two, and their thoughts on the robots’ effectiveness. This phase aimed to further explore their perceptions and gather insights on potential improvements. The study took about 50 minutes, and participants were compensated 15 USD. This study was approved by our institute’s IRB.
@@ 6.3. Participants
We recruited 10 participants from our university. 2 participants were removed due to significant technical issues. The remaining 8 participants aged from 24 to 30 years old, including 7 males and 1 female. All participants reported a high familiarity with both robots ($M=5.0$, $SD$ = 1.93) and Large Language Models ($M=6.15$, $SD$ = 1.26) on a 7-point scale, where 1 indicated ”Not at all familiar” and 7 indicated ”Very familiar.” Additionally, participants reported high frequency of LLM use, with an average score of 6.1 (SD = 1.25) on a 7-point scale, where 1 represented ”Never” and 7 represented ”Daily.” For 4 scenarios out of 32 scenarios, a different safety violation was detected: a lack of proper grounding of electrical equipment, a yellow chair on the floor, a metal rack that is not properly secured to the wall and a table that is not properly secured to the wall. Since the study has a small sample size and a re-enactment of real-world scenarios, strong conclusions should not be drawn from our results.
@@ 6.4. Results
Due to the small sample size, we performed the statistical test using the non-parametric Friedman Test. As Friedman Test only allows one variable, we reorganized the data into four levels (M-CoDAL-Low, M-CoDAL-High, GPT-Low, GPT-High)
Persuasiveness
A Friedman Test reveals there was a significant difference between the perceived persuasiveness ($X^{2}$= 15.972, p = 0.001). A Conover’s post hoc comparison showed that both M-CoDAL conditions were rated significantly more persuasive than GPT-Low (¡ .001, ¡ .001) and GPT-High (0.006 for M-CoDAL-Low and 0.021 for M-CoDAL-High). The comparison shows that M-CoDAL was rated more persuasive than GPT in all situations.
Competence
We found no significant difference in the robot’s perceived competence ($X^{2}$= 6.154, p = 0.104). Overall, participants rated the robot competent across all conditions (M = 5.92 SD = 1.98).
Discomfort
@@
Participants expressed that they perceived M-CoDAL as more intelligent and knowledgeable, with some specifically highlighting its ability to anticipate risks. For instance, P1 remarked, “The second robot (M-CoDAL) was arguing about how the twisted wires could be a potential hazard, even though I was not aware of it.” This participant had believed that placing the wire on the table was safe but appreciated how the robot made them aware of potential problems. Similarly, P10 mentioned, “I tried to convince it that keeping the cables on the table won’t be a hazard. But it told me if you knock the cables off, then it can become a hazard.” The participant noted M-CoDAL’s capacity to gather information and estimate future risks as an indicator of its intelligence.
@@ 7. Conclusion and Future Work
In conclusion, this work presents a novel approach for embodied agents to detect and engage users through dialogue when the agents detect unsafe scenarios. By leveraging coherence relations, our proposed system M-CoDAL interprets and responds to safety violations in multimodal dialogues more effectively. We introduce a novel method for active learning in generative setting for embodied conversational agents. By using an external large language model (LLM) to assess the informativeness of instances—based on safety, resolution, and user sentiment-across different clusters, we achieve broader coverage of safety scenarios, reflected in higher performance in automated evaluation. Our real-world user study demonstrates that robots equipped with M-CoDAL are viewed as more persuasive when addressing safety-related situations. This underscores the system’s potential effectiveness in real-world environments. While promising, users highlighted the agent could be irritating and annoying, highlighting the need for future systems to further personalize the coherence relations based on user response. The algorithm presented in this work, along with the publicly available multimodal dataset, offers a foundation for future studies on the deployment of proactive safety-aware embodied agents.
@@
We would like to thank Anthony Sicilia for his valuable feedback. We would like to extend our thanks to the annotators and study participants for their valuable time and effort. This work is partially funded by the National Science Foundation (Grant IIS-2112633)